Maria is 6 yrs old
Ted is one year older than Maria
Ann is 2 yrs older than Ted
How old is Ann ?

Answer by lenny460(1073)

Maria is 6 yrs old
Ted is one year older than Maria
Ann is 2 yrs older than Ted
How old is Ann ?

Maria is 6 years old
Ted = 1 + 6 = 7 years old
Ann = 2 + 7 = 9 years old

Answer: 9 years old



lennox obuong
algebra tutor